# Immersed figure-8 annuli and anyons

### v2

@@ B.2 Automorphisms utilizing tunneling
Let $\mathcal{W}$ be a punctured torus immersed in a disk $B^{2}$, as shown in the left figure of Fig. 15. If there is an Abelian sector on figure-8 region, $\mu\in{\mathcal{C}}_{8}$, then we can design an automorphism of $\Sigma(\mathcal{W})$ utilizing the tunneling trick in one of the steps; see the process shown in Fig. 15. The entire process takes the immersed punctured torus $\mathcal{W}$ back to itself. Steps 1 and 2 are smooth deformations. [^19] As a side note, I described steps 1 and 2 in a math student seminar in 2020 [41]; see the first backup slide, which was my attempt to ask about an unknown-to-me terminology. Later, my collaborators and I learned in the projects [16, 17] that the term I looked for was immersion and regular homotopy. Now step 3 is added, which gives a reason to ask more questions. Step 3 is a tunneling. Thus, the automorphism cannot be associated with a closed loop in $\mathfrak{m}(\mathcal{W},B^{2})$. Is there any physical significance to this process? We anticipate that the answer is yes, relating this to the problem of defining topological spins from a state, but we postpone the discussion to future work.
@@ Appendix C Strong isomorphism conjecture in arbitrary dimensions
Besides, the strong isomorphism conjecture (Conjecture 14) seems to resonate with another basic question: “What is the definition of quantum dimension for an arbitrary immersed sectorizable region $S\looparrowright\mathbf{S}^{n}$?” This question is open for dimensions $n\geq 3$. Many (if not all) nontrivially immersed $S$ have an embedded version $S^{\star}\hookrightarrow\mathbf{S}^{n}$, where $S^{\star}$ is in the homeomorphism class of $S$. Since $S^{\star}$ is embedded, the quantum dimension of a superselection sector $I\in{\mathcal{C}}_{S^{\star}}$ (characterized by $\Sigma(S^{\star})$) can be defined from the entropy difference of an extreme point $\rho^{I}$ with the vacuum $\sigma$:

### v4

@@ B.2 Automorphisms utilizing tunneling
Let $\mathcal{W}$ be a punctured torus immersed in a disk $B^{2}$, as shown in the left figure of Fig. 18. If there is an Abelian sector on figure-8 region, $\mu\in{\mathcal{C}}_{8}$, then we can design an automorphism of $\Sigma(\mathcal{W})$ utilizing the tunneling trick in one of the steps; see the process shown in Fig. 18. The entire process takes the immersed punctured torus $\mathcal{W}$ back to itself. Steps 1 and 2 are smooth deformations. [^22] As a side note, I described steps 1 and 2 in a math student seminar in 2020 [45]; see the first backup slide, which was my attempt to ask about an unknown-to-me terminology. Later, my collaborators and I learned in the projects [16, 17] that the terms I looked for were immersion and regular homotopy. Now step 3 is added, which gives a reason to ask more questions. Step 3 is a tunneling process. Thus, the automorphism cannot be associated with a closed loop in $\mathfrak{m}(\mathcal{W},B^{2})$.
Can we make use of the topological process in Fig. 18 to understand nontrivial universal properties of gapped topologically ordered phases? We anticipate that the answer is yes. An idea is to relate this to the problem of defining and extracting topological spins from a single wave function; we postpone the discussion to future work.
@@ Appendix C Strong isomorphism conjecture in arbitrary dimensions
Besides, the strong isomorphism conjecture (Conjecture 17) seems to resonate with another basic question: “What is the definition of the quantum dimension for a superselection sector associated with an arbitrary immersed sectorizable region $S\looparrowright\mathbf{S}^{n}$?” For dimensions $n\geq 3$, this question is open. Many (if not all) nontrivially immersed $S$ have an embedded version $S^{\star}\hookrightarrow\mathbf{S}^{n}$, where $S^{\star}$ is in the homeomorphism class of $S$. Since $S^{\star}$ is embedded, the quantum dimension of a superselection sector $I\in{\mathcal{C}}_{S^{\star}}$ (characterized by $\Sigma(S^{\star})$) can be defined from the entropy difference of an extreme point $\rho^{I}$ with the vacuum $\sigma$: